OpenStudy (anonymous):

Is my answer correct? :) Simplify the expression 1/4^-4 Answer: -1/256

OpenStudy (firejay5):

it's 256

OpenStudy (anonymous):

what?? how?? hows it not a negative????

OpenStudy (firejay5):

1/4 \[\frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

four to the fourth power equals 256, and that's without the negative sign

OpenStudy (phi):

-4 exponent means you can "flip" and change the exponent to +4 if you had \[ 4^{-4} = \frac{1}{4^4} \] or \[ \frac{1}{4^{-4}} =4^{4} \]

OpenStudy (anonymous):

So like the reciprocal?

OpenStudy (phi):

yes, an negative exponent means make it a reciprocal if you want a positive exponent

OpenStudy (anonymous):

that makes sense :) thank you for your help!!!!! :))

OpenStudy (firejay5):

was I right

OpenStudy (phi):